OpenStudy (anonymous):

CHECK MY 5 ANSWERS HELP I BEG OF YOU. PLEASE :C A crater shaped like a cone has a height of 24 yd and a diameter of 72 yd. What is the volume of the crater? Use 3.14 to approximate pi and express your final answer in hundredths. ____yd3 (i got 3000) A package of modeling clay was shaped like a cone with a radius of 24 cm and a height of 6 cm. Miriam used all the clay to make a cylinder with a radius of 16 cm. What was the height of the cylinder? Use 3.14 to approximate pi, and express your final answer in tenths. ___cm (i got 225)

OpenStudy (anonymous):

Dosha is creating a new dessert that has two layers shaped like cones. The inner cone is frozen ice cream and has a diameter of 12 cm and a height of 6 cm. The outer layer is a thin wafer shell, like an upside-down ice cream cone, with a height of 15 cm and the same diameter as the inner layer. Dosha will inject a cream filling into the space. What is the volume of the cream filling? Use 3.14 to approximate pi and express your final answer in hundredths. ___cm3 (i got 837.3 cm3)

OpenStudy (anonymous):

Volume of a Cone = (1/3)*(pi)(radius)^2 *(height) Radius = (diameter)/2 = 72/2 = 36 pi = 3.14 height = 24 so Volume of this Cone = (1/3)*(3.14)*(36)^2 *(24) = 32555.52

OpenStudy (anonymous):

for the first one

OpenStudy (mathstudent55):

Let's do the first one. Do you know the formula for the volume of a cone?

OpenStudy (anonymous):

V=πr2h 3

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

Correct. \(\Large V_{cone} = \dfrac{1}{3} \pi r^2 h\) Here, we have d = 72 yd, so r = 36 yd. h = 24 yd.

OpenStudy (mathstudent55):

\(\Large V_{cone} = \dfrac{1}{3} (3.14)(36~yd)^2 (24~yd)\) What number do you get for the volume?

OpenStudy (anonymous):

32555.52

OpenStudy (anonymous):

is it right @mathstudent55

OpenStudy (mathstudent55):

Correct.

OpenStudy (mathstudent55):

V = 32555.52 yd^3

OpenStudy (anonymous):

but it says express your final answer in hundredths

OpenStudy (mathstudent55):

What does that mean? Does it mean to round off to the nearest hundredth? If so, it's done.

OpenStudy (anonymous):

oh okay then, thanks mate :D

OpenStudy (mathstudent55):

yw